OpenStudy (anonymous):

use the euclidean algorithm to find the greatest common divisor of 524 and 467. How many divisions are required?

OpenStudy (anonymous):

To do the euclidean algorithm for 201 and 43 \[ 201 = (4)(43) + 29\] \[43 = (1)29 + 14\] \[29 = 2(14) +1\] \[14 = (14)(1) +0\] Taking the remainder before 0, the gcd of 201 and 43 = 1. or (201,43)=1